OpenStudy (anonymous):

The terminal speed (in m/s) of the skydriver can be approximated by square root 40 m, where m is the mass (in kg) of the skydiver. calculate the terminal speed of a 100-kg shydiver. write an exact answer in radical form. simplify the radical by factoring.

OpenStudy (bahrom7893):

Okay let's try to solve this in simple steps.

OpenStudy (bahrom7893):

Terminal velocity equation: \[V_t = \sqrt((2mg)/(\rho*A*C_d))\]

OpenStudy (bahrom7893):

(everything is under square root)

OpenStudy (bahrom7893):

where: \[V_t\] is terminal velocity m is mass of the falling object g is acceleration due to gravity (9.81m/s^2) Cd = drag coefficient ρ = density of the fluid through which the object is falling A = projected area of the object.

OpenStudy (bahrom7893):

In your case, the terminal Velocity is \[\sqrt{40*m}\] where m is the mass (in kg) of a skydiver.

OpenStudy (bahrom7893):

So if you need the terminal velocity of a skydiver whose mass is 100kg, then plug in 100kg into the equation above... You didn't even need the terminal velocity equation and all that crap.. sorry I didn't read the question carefully first time

OpenStudy (bahrom7893):

In your case, the terminal Velocity is \[\sqrt{40*m}\] where m is the mass (in kg) of a skydiver. If the skydiver's mass is 100kg, then the terminal velocity is: sqrt( 40 * 100 ) = sqrt(400 * 10) = 20sqrt(10)

OpenStudy (bahrom7893):

The answer is 20 * sqrt(10) m/s

OpenStudy (anonymous):

wow that helped alot thanks